### v1

# LongVILA: Scaling Long-Context Visual Language Models for Long Videos

Abstract
Long-context capability is critical for multi-modal foundation models. We introduce LongVILA, a full-stack solution for long-context vision-language models, including system, model training, and dataset development. On the system side, we introduce the first Multi-Modal Sequence Parallelism (MM-SP) system that enables long-context training and inference, enabling 2M context length training on 256 GPUs. MM-SP is also efficient, being 2.1×2.1$\times$2.1 × - 5.7×5.7$\times$5.7 × faster than Ring-Style Sequence Parallelism and $\times1$ 1.4×1.4$\times$1.4 × faster than Megatron-LM in text-only settings. Moreover, it seamlessly integrates with Hugging Face Transformers. For model training, we propose a five-stage pipeline comprising alignment, pre-training, context extension, and long-short joint supervised fine-tuning. Regarding datasets, we meticulously construct large-scale visual language pre-training datasets and long video instruction-following datasets to support our multi-stage training process. The full-stack solution extends the feasible frame number of VILA by a factor of 128 (from 8 to 1024 frames) and improves long video captioning score from 2.00 to 3.26 (1.6x), achieving 99.5% accuracy in 1400-frames video (274k context length) “needle in a haystack”. LongVILA-8B also demonstrates a consistent improvement in performance on long videos within the VideoMME benchmark as the video frames increase.
1 Introduction
The importance of Multi-Modal and Long Context From using customized machine learning models to Artificial General Intelligence (AGI), AI research is moving towards a unified solution for everything. To this end, the most exciting breakthrough of foundation model research is unlocking new abilities (Wei et al., 2022), which indicates that the model offers new potential to solve both existing challenges and unforeseen applications. Two important directions of unlocking abilities are multi-modality and long-context.
First, a foundation model supporting more modalities can take more flexible input signals so that people can interact with the model in more diverse manners, e.g., GPT-4o-like multi-modal chatbot, multi-modal web agent (Koh et al., 2024), and real-world robotics foundation model (Brohan et al., 2022; 2023; Padalkar et al., 2023). The longer context enables models to process more information, e.g., long documents, repo-level codebase, and hour-length video, which similarly provides required features to more real-world applications. Since we expect the foundation model to work as a general solution for everything, supporting these two features is highly desirable to achieve real AGI.
Recently, there are many works that enable VLM for short context(Lin et al., 2023b), as well as LLM for long context(Chen et al., 2023; Li et al., 2023a). However, there aren’t much work focusing on Multi-Modal + Long Context at the same time. LWM(Liu et al., 2024a) proposes Ring Attention to extend context length, however the discrete tokens in vision part lead to low performance of image/video understanding although it simplifies the system design. LongVA(Zhang et al., 2024a) proposes training long context LLM and vision encoder on short videos and extend context window during inference. It can help increase context window for VLM but hits limit.
The Necessity of Full-Stack Design in Long Multi-Modal Language Model Compared to the traditional machine learning algorithms (Jordan & Mitchell, 2015) Large Language Model (LLM) training is usually a more systematic project requiring more data engineering (Betker et al., 2023; Ouyang et al., 2022; Zhou et al., 2024) and hardware-software co-design (Lepikhin et al., 2020; Chowdhery et al., 2023; Shoeybi et al., 2019; Brown et al., 2020; Dehghani et al., 2023). Moreover, compared with text-only LLM, the multi-modal foundation model and long-context model both introduce more challenges requiring the multi-component co-design: (1) multi-modal foundation models (e.g., LLaVA (Liu et al., 2023b), and Flamingo (Alayrac et al., 2022)) usually have different model architectures with the text-only LLMs, so that we might need a different distributed training strategy to support the large multi-modal transformer architecture; (2) long context modeling requires both long context data to unlock the ability (Fu et al., 2024b; Chen et al., 2023) and also the infrastructure to make memory-intense long context training possible (Li et al., 2021; Jacobs et al., 2023; Li et al., 2023b). As a result, if we want to introduce both long-context and multi-modal abilities into the foundation model, a careful full-stack design is required to scale the model size and context length up together.
Figure 1: The training pipeline for the LongVILA model extends existing VLM models to support long video processing. Using VILA-1.5 (Lin et al., 2023b) as an example, the process begins with alignment, pre-training, and supervised fine-tuning, in Stages 1 through 3. In Stage 4, the model undergoes mid-training for context extension using the VILA model. Finally, in Stage 5, the model is fine-tuned for long video understanding with MM-specific supervised pretraining (MM-SP), resulting in the LongVILA model. While this example uses VILA, the process is adaptable to other VLMs. Note that we train LongVILA models for up to 1024 video frames in Stage 5.
LongVILA: All-in-One Long-Context Visual Language Model Scaling In this work, we devote ourselves to an all-in-one solution for the training and inference of long-context visual language model. In training, we need to collect (1) a large amount of high-quality visual language data to conduct large-scale pre-training; and (2) high-quality long-context visual language instruction following data to unlock the long-range modeling ability. In addition to the data, we must implement the efficient and easy-to-use infrastructure, i.e., Multi-Modal Sequence Parallelism (MM-SP), supporting (1) multi-modal LLM architecture and (2) memory-intense long context transformer training. From the modeling perspective, we conduct a five-stage training curriculum: (1) multi-modal alignment, (2) large-scale pre-training, (3) short supervised fine-tuning, (4) context extension for LLMs, and (5) long supervised fine-tuning. For inference, the memory usage of KV cache will also be a bottleneck when the sequence length is very long, we thus implement the inference mode of our MM-SP to support long context multi-modal language deployment.
Contributions The contributions of our work can be summarized as follows:
  • From the system level, we proposed Multi-Modal Sequence Parallelism(MM-SP), the first open-sourced sequence parallelism implementation supporting the training and inference of long-context multi-modal language models. Our system is efficient and extremely easy to use (i.e., natively compatible with Huggingface Transformers [^1] https://huggingface.co/docs/transformers/en/index).
  • From the data level, we took both large-scale visual pre-training and long-context multi-modal scaling into consideration. Our data receipt enables our model to generalize well on various domains with 256k tokens.
  • We developed a working training recipe to turn a short context LLM and vision encoder into a long context VLM.
  • As a result, our model achieved competitive results on both video caption and video QA tasks at varied context lengths, which outperforms state-of-the-art open-sourced models by a large margin.
2 Background and Related work
Figure 2: Scaling the context length can improve multi-modality accuracy, including both Video-QA on VideoMME and Video Captioning on LongVILA-Caption.
Visual language model architecture For VLMs, there are two mainstream designs, i.e., encoder-decoder VLM (e.g., LLaVA (Liu et al., 2023b), PaLM-E (Driess et al., 2023)) and decoder-only VLM (e.g., Fuyu (Bavishi et al., 2023), Chameleon (Team, 2024)).
  • Decoder-only VLMs still lag behind the encoder-decoder VLMs although there are some positive signals showing the potential of decoder-only VLMs. There might be various reasons behind this results.
  • Encoder-Decoder VLMs connect the vision encoder and LLM decoder via the multi-modal projector. Some of the multi-modal projectors (e.g., spatial pooling, Q-former) can significantly reduce the number of tokens for each image or video frame, so that greatly reduce the computational cost of the LLM decoder. However, decoder-only LLMs take raw patches as input without hierarchical token pooling by default, so that make it more challenging to reduce the number of tokens for each image or frame. In this work, we follow VILA (Lin et al., 2023b) as the starting point. Note that there are some other enhanced variants of VILA, including VILA2 (Fang et al., 2024) for better performance and the X-VILA (Ye et al., 2024) for cross-modality understanding, reasoning, and generation. We follow the plain VILA-1.5 version for model architecture and training pipeline.
Sequence parallelism and hybrid strategy Long-context training examples typically produce activation beyond the available memory of one device. To tackle this memory issue, multiple systems have employed the sequence parallelism paradigm in text-only LLMs community, which distributes a single sequence into multiple devices. In particular, Ring-style systems Li et al. (2021; 2023b); Liu et al. (2023a) uses Point-to-Point (P2P) communication primitives to collectively compute the attention module, while DeepSpeed-Ulysses Jacobs et al. (2023) uses All-to-All (A2A) primitive to change the sharding dimension between sequence dimension and attention head dimension to compute the attention module. Our paper is the first to design and implement a sequence parallelism system for visual language models. Furthermore, we carefully study the A2A and P2P primitives and design a novel hybrid sequence parallelism strategy to optimize training throughput. (Gu et al., 2024) is a concurrent work which also investigates the integration of these two SP strategies. However, it is limited to vanilla LLM training and does not extend to multi-modal scenarios due to the lack of support for complex attention mask inputs and variable-length input sequences.
3 System
Long-context VLMs training induces excessive memory footprints. For instance, during our Stage 4 long video training (Figure  1), a single sequence consists of 128K tokens, whose activation consume more memory than available in a single GPU. The most popular open-sourced solution, fully sharded data parallelism does not distribute the activation produced by a single sequence, and thus can not support our use case. Thus, we develop our own system based on sequence parallelism (Li et al., 2021; 2023b; Liu et al., 2023a; Jacobs et al., 2023), commonly implemented in existing foundation model system to optimize text-only LLM training. However, we find that existing systems are not efficient and not scalable enough to support our long-context VLMs workload.
Figure 3: MM-SP comprises of our novel 2D attention technique that uses intra-node All2All and inter-node P2P to communicate keys and values (KV), and a novel sharding strategy that balances the computation of both the image encoder and the language modeling stage. In contrast, Ring-style SP (Liu et al., 2023a; Li et al., 2021; 2023b) employs P2P in both an intra-node and an inter-node setting, causing excessive communication overhead. In addition, they are developed in text-only modality, where the workload of an image encoder is not optimized. Intra-node and inter-node bandwidths are measured in internal H100 clusters.
3.1 Limitations of Existing Systems
In this section, we discuss the inefficiency of current sequence-parallel systems when applying multi-modal LLMs.
Modality heterogeneity In text-only LLMs, sequences are processed by a single tokenizer into tokens, where it is straightforward to distribute tokens evenly into multiple GPUs. However, VLMs include an encoder architecture, where a non-text data is usually first represented by a single place-holder token (e.g. $<$img$>$), and encoded with various real tokens during training. For example, a video frame usually take around 256 tokens (Lin et al., 2023b). Due to the different processing in visual and text modality, a simple implementation that treats place-holder tokens equally as a text token results in a workload imbalance between GPUs (Figure  3).
Network heterogeneity Our multi-modality comprises extremely long videos (Figure 1), which requires employing sequence parallelism in a multi-node setting. In a multi-node setting, inter-node and intra-node network bandwidth differs significantly. For example, the NVIDIA DGX H100 utilizes NVLink at 900 GB/s for intra-node GPU communication and InfiniBand at 50 GB/s for inter-node GPU communication (single path), resulting in an 18$\times$ difference in bandwidth. Previous work, Ring-Style sequence parallelism (Li et al., 2021; 2023b; Liu et al., 2023a; Zhu, 2023) ignores the heterogeneous networking feature on GPUs and utilizes P2P communication in both inter-node and intra-node settings. This design induces excessive communication costs where they usually attempt to overlap them into computation. However, we found that this design cannot always hide the overhead, and even slows down the computation kernel (Table 1).
Table 1: The forward and backward attention kernel wall-clock time with or without the overlapping design (Unit: $\mu s$). The communication overlap designs in Ring-style SP slows down the attention kernel by acquiring streaming multiprocessor (SM).
Seq. length 4K 8K 16 K 24K 32K
forw. w/o 29.5 49.3 122.1 239.2 402.9
forw. w/ 35.0 (+18.6%) 54.6 (+10.7%) 131.2 (+7.5%) 250.9 (+4.8%) 420.1 (+4.2%)
backw. w/o 77.7 123.3 362.9 730.0 1218.9
backw. w/ 82.2 (+5.8%) 129.8 (+5.3%) 367.0 (+1.1%) 743.2 (+1.8%) 1225.3 (+0.5%)
Limited maximal sequence length supported DeepSpeed-Ulysses (Jacobs et al., 2023) offers a potential solution to address the communication problem in ring-style SP. It uses All-to-All communication primitives, which induces less communication volumn. Unfortunately, this use of All-to-All primitives come with costs - its design is bind to paralleling in the attention head dimension, instead of the sequence dimension in the attention computation. Thus, simply using DeepSpeed-Ulysses cannot scale beyond the number of attention heads. For example, Llama-3 8B model employs GQA with 8 Key-Value heads, limiting the maximum sequence parallelism degree to 8. Even further employing the replication approach for Key-Value heads with additional communication overhead (Li et al., 2023b), the highest achievable sequence parallelism degree is still restricted to 32 (number of Query heads). This is not sufficient when handling extremely long sequences (e.g, full movies).
3.2 Multi-Modal Sequence Parallelism Training Mode
Observing problems in existing systems, we summarize that an ideal multi-modal sequence parallelism should achieve efficiency and scalability, by addressing the modality heterogeneity and the network heterogeneity, and also be able to scale beyond the number of attention heads.
Figure 4: Workflow of Multi-Modal Sequence Parallelism, where Batch Size (bs) = 4 and Sequence Parallel Size (SP_Size) = 3. To enable sequence parallelism for multi-modal inputs, we implement a tailored sharding strategy that achieves balanced workload distribution and is compatible with our 2D-attention mechanism.
MM-SP workflow To address the issue of modality heterogeneity, we propose a two-stage sharding strategy that optimizes the compute workload for both the image encoding and language modeling stages. As depicted in Figure 4, the process begins with evenly distributing images (e.g., video frames) across devices within the SP process group, thereby achieving load balancing during the image encoding stage. Subsequently, we gather the global vision and text inputs for token-level sharding in the second stage. To support ring-based attention, we extend sequences with arbitrary dummy tokens, ensuring each sequence can be divided by the ring-based SP degree. This modification maintains consistency with the original version by adjusting label inputs to ignore the padded tokens when calculating loss. We employ a balanced sharding strategy to distribute the context to each rank from both ends, ensuring equal computation across ranks. The advantage of this strategy will be demonstrated later (Table 4). Since this redistribution is performed only once during training, the overhead is minimal. Finally, the balanced local inputs are fed into the LLM backbone, utilizing the 2D-attention mechanism to achieve efficient sequence parallelism.
Figure 5: Comparison of RingAttention, ZigZag-RingAttn, and 2D-Attention. The blue block represents there is commutation between QKV and the black frame indicates local attention computation on each SP group rank (SP_Size=3).
2D-Attention mechanism To address the network heterogeneity and achieve scalability, we leverage the advantages in ring-style SP and Ulysses SP. More specifically, we treat paralleling across the sequence dimension or the attention head dimension as a “1D SP”. Our method parallelizes computation across both attention head and sequence dimension. In particular, we convert 1D SP process group into a 2D mesh, including independent Ring (P2P) process groups and Ulysses (A2A) process groups. For example, as shown in Figure 3 left, to enable an 8-degree sequence parallelism across 2 nodes, we can construct a 4$\times$2 communication mesh using 2D-SP. In this configuration, the A2A process group, with a size of 4, distributes the QKV tensors according to the head dimension and re-partitions them along the sequence dimension within each node. Meanwhile, the P2P process group, with a size of 2, transfers the partitioned KV chunks between nodes. Besides, to further illustrate how ZigZag-RingAttn balances computation and how 2D-attention works, we depict the attention computation schedule using different methods in Figure 5. In this example, the sequence length is 8 and the global SP degree is 4. Due to the triangular structure of causal attention computations, RingAttention experiences a computation imbalance, with rank 0 becoming idle after the first round while rank 3 continues computing through all stages. ZigZag-RingAttn  addresses this by reordering input sequence tokens along the sequence dimension to achieve load balance. Our 2D-attention mechanism uses a ring parallel degree of 2 and a head parallel degree of 2, resulting in an equivalent global sequence parallel degree of 4. This approach also incorporates a workload balance strategy within the ring-based process group. Additionally, it employs the All2All operation to distribute the QKV tensors based on the head dimension across devices, ensuring efficient and balanced computation.
3.3 Multi-Modal Sequence Parallelism Inference Mode
The model we obtained by sequence parallelism training can handle long-context multi-modality downstream tasks. However, the most widely used inference system built on HuggingFace Transformers is usually ran on a single GPU. A lack of a distributed implementation hinders the maximum sequence length we can perform inference. The most straightforward and easy-to-use solution in HuggingFace Transformers is to use its own pipeline parallelism inference feature, which shards a single model across different devices in a layer-basis (Huang et al., 2019; Narayanan et al., 2019). However, this approach is inefficient, where only one GPU is activated at the same time. It also does not well support long sequence length because the first device needs to store more big tensors such as the inputs embedding and images, bottlenecking the whole system.
Figure 6: GPU schedule of HuggingFace pipeline versus MM-SP inference mode, demonstrated with 4 GPUs. MM-SP leverages all GPUs to jointly compute for inference results.
Thus, we implement sequence parallelism for VLMs distributed inference. Compared to the training mode, the system needs to additionally maintain tensors (e.g. input tokens and position encodings) that are progressively changing during the decoding phrase (Yu et al., 2022). In addition, the system needs to detect signals from the machine that holds the last token and accordingly terminate the distributed process. Compared to the Huggingface native pipeline parallelism strategy, our inference mode is efficient in that all devices jointly compute during the inference, accelerating the process by the number of machines times (Figure 6). In addition, it is scalable in that all memory is evenly distributed, and thus can support longer sequences with more machines.
4 Training and Data
Figure 7: The pipeline for generating instruction-following data from long videos. The process begins by segmenting the long video into short clips, each approximately 10 seconds in length. These clips are individually annotated with captions using the VILA-1.5 model. Subsequently, a Large Language Model (LLM) is employed to generate question-and-answer pairs based on the captions of these clips. The types of questions generated include summarization and other inquiries pertinent to the content of long videos.
Figure 8: The proportion of question and video categories in our LongVILA__\__sft dataset. We have 15,292 videos in total. For each video, there are one sample for captioning and the other question. The left pie chart shows the proportion of video categories. The right pie chart shows the proportion of question categories.
4.1 Overall Training Pipeline
As shown in Figure 1, in our pipeline, there are five training stages, i.e., Stage 1: multi-modal alignment, Stage 2: large-scale pre-training, Stage 3: context extension for LLM, Stage 4: supervised fine-tuning. Stage 1 and Stage 2 are following VILA 1.0 Lin et al. (2023b), to firstly bridge the gap between LLM and vision encoder, and then pre-training on larger datasets. In Stage 1, only the multi-modal projector is trainable with others frozen. In Stage 2, We frozen the vision encoder and training LLM and the multi-modal projector. Afterwards, we exend the context length of LLM with text-only dataset in a continued pre-training manner. In Stage 4, we enhance the instruction following abilities by multi-task supervised fine-tuning. Both short and long data, including images, short videos and long videos related questions, are used for training. It is noted that all parameters are trainable in the final stage.
4.2 Stage1&2&3: Multimodal Alignment, Pre-training, and Short Supervised Fine-tuning
Before training at scale, we first use open-sourced image and video caption datasets to train the multi-modal projector in stage (1) to conduct the multi-modal alignment. Note that, following (Lin et al., 2023b), both vision encoder and LLM decoder are frozen at this stage. After that, we conduct large-scale pre-training to learn general multi-modal capacity at scale. To improve the quality of large open-sourced datasets, we follow VILA2 (Fang et al., 2024) to relabel COYO-25M (Lin et al., 2023b; Byeon et al., 2022) with VILA-1.5-40B (Lin et al., 2023b). The supervised fine-tuning process incorporates mixed data types, including both images and videos. For short video comprehension, we utilize open-source video instruction-following datasets, e.g., YouCook2 Zhou et al. (2018) and ShareGPTVideo Zhang et al. (2024b).
Table 2: Training system throughput comparison against various Megatron-LM strategy on 32 H100 GPUs, measured in time per iteration (unit: second). ”OOM” stands for ”Out of Memory”.
Sequence length Megatron-LM Ours
CP CP=4+TP=8 ZigZag-RingAttn Ulysses 2D Attention
320 K OOM OOM 23.57 10.70 11.12
288 K OOM OOM 20.24 8.68 8.65
256 K OOM OOM 17.54 6.98 7.04
224 K OOM OOM 15.22 5.47 5.53
192 K OOM OOM 12.97 4.15 4.24
160 K OOM OOM 10.83 3.02 3.11
128 K OOM 2.96 8.38 2.07 2.17
96 K 4.32 1.77 6.35 1.33 1.41
64 K 3.00 0.96 4.25 0.76 0.80
32 K 1.72 0.44 2.26 0.39 0.40
4.3 Stage4: Context Extension for LLMs
Our empirical research indicates that extending the context length of LLMs is essential prior to engaging in supervised fine-tuning with long video datasets. Following Stage 2 of our methodology, we execute a continuation of pre-training on the LLM to enhance its context length to 262,144, utilizing a total of 17B tokens. We employ a progressive training schedule, incrementally increasing the context length from 8,192 to 65,536, and ultimately to 262,144, utilizing the SlimPajama dataset (Soboleva et al., 2023) in accordance with the methodology outlined by (Fu et al., 2024c).
Furthermore, we augment the base frequency of the Rotary Position Embeddings (RoPE) as described in (Su et al., 2021) during the fine-tuning phase. Sequence parallelism is implemented for the training at the 262,144 context length. We use low-rank adaptation for context extension fine-tuning (Chen et al., 2024b). These processes collectively require approximately 336 GPU hours on machines equipped with 80GB A100 GPUs.
4.4 Stage5: Long Supervised Fine-tuning
Long video instruction following dataset To facilitate the fine-tuning of long videos, we also constructed a new, dedicated dataset for long video training, consisting of 15,292 videos with around 10 minute for each one. We use the original long videos from the Shot2Story dataset (Han et al., 2023). Each video includes different questions and answers: one for generating captions and another for answering questions, enabling diverse applications in video understanding. Figure 7 illustrates the process for generating instruction-following datasets from long videos. Initially, the long video is segmented into shorter clips, each approximately 10 seconds in duration. These clips are then independently annotated with descriptive captions utilizing the VILA-1.5 model. Subsequently, an LLM is employed to generate question-and-answer pairs derived from the captions of these clips. The generated questions encompass summarization and other queries relevant to the comprehensive understanding of long video content.
As in Figure 8, the left chart categorizes videos into several domains, including Travel & Events, Sports, Education, Pets & Animals, People & Blogs, News & Politics, Music, Science & Technology, Comedy, Entertainment, Film, and Gaming, ensuring a wide-ranging representation of video content. The right chart breaks down the categories of questions into Spatial, Attribute, Action, Object, OCR, Synopsis, and Temporal, reflecting the variety of inquiries and cognitive tasks that the dataset can address. This dataset provides a rich resource for advancing the understanding and processing of long video formats in supervised fine-tuning.
5 Evaluation
We evaluate our full-stack solution from both system and modeling perspectives. We report our training and inference results first to verify the efficiency and scalability of our system, which makes long-context training and inference possible. We evaluate the captioning and instruction following of our long-context model.
5.1 Training and inference system
Figure 9: Maximal supported training sequence length comparison against ZigZag-RingAttn (Li et al., 2023b; Liu et al., 2024a; Zhu, 2023; Korthikanti et al., 2023) and DeepSpeed Ulysses (Jacobs et al., 2023). Our system supports higher training sequence length compared to Ulysses, and on par with ZigZag-RingAttn.
Our developed training and inference systems can be used with HuggingFace Transformer by simple monkey patching, following the popular open-source convention in (Zheng et al., 2023). In this section, we provide quantitative evaluation of training system throughput, inference system latency, and the maximal supported sequence length.
5.1.1 Training system
Baselines and hardware setup For training efficiency, we compare with a ZigZag ring-style sequence parallelism system, which is incorporates load-balanced and GPU optimization (ZigZag-RingAttn for consistency) (Li et al., 2023b; Zhu, 2023; Liu et al., 2024a; Korthikanti et al., 2023). We select a popular open-sourced implementation (Zhu, 2023). We use Fully-sharded-Data-Parallelism (FSDP) (Zhao et al., 2023) to reduce the memory occupied by models, gradients and optimizer states instead of Zero-3 (Rajbhandari et al., 2020) (Table  9). In addition, we compare with the expert-designed and highly optimized Megatron-LM (Shoeybi et al., 2019; Korthikanti et al., 2023) system. We compare with their implementation of sequence parallelism, termed “context-parallelism” (CP). We also compare with a hybrid strategy that uses tensor model parallelism (TP) within a node, and CP cross node, recommended by the Megatron-LM team as an advanced usage.
We conduct most experiments on H100 nodes, each with 8xH100 (80GB) machines equipped with intra-node NVLink, and 400 Gbps inter-node Infiniband. We conduct the maximal supported sequence length supported training experiments up to 32 A100 nodes, each equipped with 8xA100 (80GB), where the same conclusion apply for H100 because of the same total memory. We evaluate on a model of size 8B and a batch size of 1. Since Megatron-LM baseline does not natively support VLMs training and that the visual encoder is usually orders of magnitude smaller than LLMs, we provide main results for the LLM backbone, without the visual encoder. We provide an ablation of the visual encoder part in § 5.1.3.
Figure 10: Training and inference system performance obtained on 32 and 8 H100s.
[…]
Table 3: First token latency (left; unit: second) and maximal supported inference sequence length (right) comparison versus native HuggingFace pipeline.
Seq. 8K 16K 32K 64K 96K 128K 256K
HF 1.81 2.47 4.20 8.50 15.36 OOM OOM
Ours 0.22 0.30 0.53 1.08 1.88 2.89 9.24
# GPUs 1 2 4 8
HF 48K 83K 96K 96K
Ours 48K 88K 138K 276K
Table 4: Time per iteration on the ShareGPT4Video dataset (Chen et al., 2024a) with and without our two-stage sharding strategy, measured with different settings as detailed in §$\S$ 5.1.3 (Unit: seconds).
2 GPUs 4 GPUs 8 GPUs
one-stage 0.78 0.89 1.20
two-stage (our MM-SP) 0.77 0.86 1.12
Figure 11: Comparison of Needle in the Long Video Haystack Experiment. The 32-frame baseline model (left) can not retrieve right needles after 32 frames. In contrast, the LongVILA model (right), trained on 1024 frames, presents 99.5% accuracy on 274k context length.
Table 5: Performance evaluation on Video MME (Fu et al., 2024a) benchmark.
Model Frames Short Medium Long Overall
Qwen-VL-Chat 4 46.9 38.7 37.8 41.1
Video-LLaVA (Lin et al., 2023a) 8 45.3 38.0 36.2 39.9
ST-LLM (Liu et al., 2024b) 64 45.7 36.8 31.3 37.9
VideoChat2-Mistral (Li et al., 2023c) 16 48.3 37.0 33.2 39.5
Chat-UniVi-V1.5 (Jin et al., 2023) 64 45.7 40.3 35.8 40.6
VideoLLaMA2 (Cheng et al., 2024) 16 56.0 45.4 42.1 47.9
LLaVA-NeXT-Qwen2 32 58.0 47.0 43.4 49.5
LongVILA-8B 128 60.2 48.2 38.8 49.2
256 61.8 49.7 39.7 50.5
In summary, our training system takes the advantage of both worlds - it achieves similar scalability as ZigZag-RingAttn, and maintain the throughput of DeepSpeed-Ulysses. It achieves 1.3x speedup and a 2.5x longer context length support compared to the highly optimized Megatron-LM.
5.1.2 Inference system
We evaluate our inference system against HuggingFace Pipeline parallelism on a single node of 8 H100 GPUs and the 8B model. (Table  3). We achieve 8.2×8.2$\times$8.2 × speed up compared to HuggingFace pipeline on 8xH100 GPUs. This is mainly because Huggingface pipeline inference only activates 1 GPU at a time, where our method leverage all GPUs to jointly compute for results. Table 3 compares the maximal sequence length supported, where we find our method achieves 2.9×2.9$\times$2.9 × longer sequence length than HuggingFace pipeline. In particular, we find that during 96K inference, the largest sequence length it can support, HF pipeline stores 80GB activations on the first GPU and 18GB on later GPUs. This imbalance allocation of activation bottlenecks the maximal sequence length.
5.1.3 Effect of two-stage sharding
We ablate the effect of our two-stage sharding strategy using a video captioning dataset (Chen et al., 2024a). We compare our two-stage sharding strategy against a baseline one-stage sharding strategy that only shard by the number of images. We measure the time per iteration under different number of H100 GPUs. For $k$ GPUs, we use a total number of $k$ images per video, and a batch size of $k$. The results are shown in Table 4. We observe a speedup from 1% to 7%. The speedup comes from the long captioning, where the baseline suffers from imbalance workload without sharing by the number of text tokens.
5.2 Needle in a Haystack
In Figure 11, we present the results of the ”Needle in a Haystack” experiment for long videos. Following the methodology established in the existing literature (Zhang et al., ), we prepared a long video and sampled a fixed number of frames. We inserted specifically designed images at various depths and tasked the model with answering corresponding questions. The 32-frame baseline model (left) was unable to accurately retrieve the correct images beyond 32 frames. In contrast, the LongVILA model (right) demonstrated enhanced performance across a range of frame counts and depths.
5.3 General Video Understanding
Table 5 presents the performance of various models on the Video MME benchmark, comparing their effectiveness across short, medium, and long video lengths, as well as overall performance. LongVILA-8B, utilizing 256 frames, achieves an overall score of 50.5. We also do the ablation on the effects of Stage 3 and Stage 4 in Table 6.
Table 6: Ablation on Stage 3: Short SFT and Stage 4: Mid-training for LongVILA-8B on Video MME benchmark with 128 frames for evaluation.
Stage 3 Stage 4 Average Short Medium Long
28.5% 29.2% 28.8% 27.4%
44.9% 55.1% 43.0% 36.5%
49.2% 60.2% 48.2% 38.8%
Table 7: Evaluation of LongVILA-Caption performance across different frame counts. LongVILA-Caption’s performance improves across all metrics. As the number of frames increases from 8 to 256, with the average score rising from 2.00 to 3.26 out of a maximum of 5. Unlike in Figure 2, the models presented in this table are both trained and evaluated using the same number of frames.
Frames Correctness Detailed Contextual Average
8 1.87 1.85 2.27 2.00
128 2.36 2.44 2.79 2.53
256 3.23 3.11 3.43 3.26
5.4 LongVILA-Caption
We have developed a long video captioning benchmark, LongVILA-Caption, consisting of 100 long videos, with captions generated as detailed in Section 4.4, and verified through human examination. In line with the methodology of VideoChatGPT (Maaz et al., 2024), we evaluate the predictions of each model based on their correctness, detailed orientation, and contextual understanding. For instance, we assess correctness by employing GPT-4 to predict scores using a specific prompt. Additionally, we present two examples in Figures 13 and 14, featuring long videos in sports and technology. These examples demonstrate that LongVILA, with its capability to process more frames, offers a more comprehensive understanding of videos compared to its short-frame counterpart.
The Table 7 presents the performance metrics for the LongVILA models being trained and evaluated on varying numbers of frames: 8, 128, and 256. As the number of frames increases, the model’s performance improves significantly. Specifically, the average scores rise from 2.00 to 3.26, highlighting the model’s enhanced capability in generating accurate and rich captions with more frames.
6 Conclusion
We present LongVILA, a full-stack solution for long-context visual language model, including distributed systems, model training, and data engineering. Our system can efficiently scale context length up to 2M tokens and achieve 2.1$\times$ - 5.7 $\times$ and 1.1$\times$ - 1.4$\times$ speedup than Ring Sequence Parallelism and Megatron-LM, respectively. We then meticulously curated long video datasets and train our model with a two-stage pipeline based on pre-trained VLM checkpoint. Our LongVILA-8B model extends the feasible frame number from 8 to 1024, capturing fine-grained information precisely from 2-hour video ”needle-in-a-haystack” and achieving promising results on both video QA and captioning.
Appendix A Appendix
Figure 12: Convergence evaluation. We compare the training loss curve of training LongVILA with and without sequence parallelism. This figure shows convergence of LongVILA-8B model on 8 A100 GPUs with sequence parallelism degree set at 4 and pure data parallelism training. The training dataset mixture includes shot2story, ai2d and chartqa. It is obvious that the two curves align closely, indicating that our MM-SP system has no adverse impact on training quality.
FSDP versus Zero-3 with hybrid sequence parallelism
Table 8: Training system throughput comparison on 64 H100 GPUs, measured in time per iteration (unit: second). Ulysses does not apply here since it can only support up to 32 GPUs.
Sequence length Megatron-LM Ours
CP CP=8+TP=8 ZigZag-RingAttn Ulysses 2D Attention
640 K OOM OOM 88.4 - OOM
578 K OOM OOM 77.2 - 16.9
512 K OOM OOM 66.1 - 13.31
448 K OOM OOM 57.5 - 10.39
384 K OOM OOM 48.6 - 7.80
320 K OOM OOM 40.5 - 5.63
256 K OOM 5.31 32.2 - 3.93
192 K 8.81 3.10 24.1 - 2.49
128 K 7.10 1.57 16.0 - 1.36
64 K 3.09 0.61 8.04 - 0.57
32 K 1.86 0.44 4.24 - 0.33
Table 9: Our training system using together with FSDP (Zhao et al., 2023) or Zero-3 (Rajbhandari et al., 2020) on 32 H100 GPUs. We found that FSDP is more friendly in memory management and choose to use it in our default configuration. (Time per iteration, unit: second).
Sequence Length Zero-3 FSDP
ZigZag-RingAttn Ulysses 2D Attention ZigZag-RingAttn Ulysses 2D Attention
320 K OOM OOM OOM 23.57 10.70 11.12
288 K OOM OOM OOM 20.24 8.68 8.65
256 K OOM OOM OOM 17.54 6.98 7.04
224 K 19.04 7.06 5.73 15.22 5.47 5.53
192 K 13.01 4.24 4.38 12.97 4.15 4.24
160 K 10.73 3.09 3.23 10.83 3.02 3.11
128 K 8.63 2.16 2.30 8.38 2.07 2.17
96 K 6.49 1.43 1.53 6.35 1.33 1.41
64 K 4.40 1.01 1.08 4.25 0.76 0.80
32 K 2.06 1.58 1.04 2.26 0.39 0.40
Figure 13: Examples of sports long video caption with LongVILA. For the gameplay opening, the 8-frame baseline describes only static image, two players in only blue jerseys. In contrast, 256-frame LongVILA describes players in blue and red jerseys passing and kicking the ball. In addition, the 256-frame version also include the detail of players hugging emphasizes the celebratory aspects, which is missing in the 8-frame baseline.
Figure 14: Examples of technology long video caption with LongVILA. At the beginning of captions, the 8-frame baseline only describes static image and two cars. In contrast, the 256-frame LongVILA describes the car on snowy road, covering front, rear, and side views. For details, the 256-frame LongVILA describes close-ups of ignition button, gear shift, and dashboard elements, which are missing in the 8-frame baseline.